Suppose a school has 960 learners enrolled. The ratio of the number of boys to girls is 4:8. Find the difference between the number of boys and the number of girls enrolled at the school.

1
Expert's answer
2022-06-14T12:09:18-0400

k - proportionality factor

Boys -4k, girls - 8k.

4k+8k=960

12k=960

k = 80

Boys: 4*80=320

Girls: 8*80=640

640-320=320 is the difference between the number of boys and the number of girls .